Prove that the coherence length is related bu bandwidths following equation:-
L=c/∆v= λ²/∆λ
1
Expert's answer
2019-08-07T09:59:56-0400

The frequency and wavelength are related through


"v=\frac{c}{\lambda} \longrightarrow \frac{dv}{d\lambda} = -\frac{c}{\lambda^2}."

Differentiating the latter we can find the frequency width "\Delta v" from the wavelength width "\Delta \lambda"

"\frac{\Delta v}{\Delta \lambda} \approx \left| \frac{dv}{d\lambda} \right|=\frac{c}{\lambda^2} \longrightarrow \Delta v=\frac{c \Delta \lambda}{\lambda^2}."

The coherence time is

"\Delta t=\frac{1}{\Delta v}."

The coherence length is

"L=c\Delta t=\frac{c}{\Delta v}=\frac{\lambda^2}{\Delta \lambda}."